Jolin had to multiply 2.85 by 3.She added the two numbers instead.What was the difference between her answer and the actual answer?
1
Expert's answer
2011-10-06T13:11:27-0400
Jolin answer
J=3+2.85=5.85
Actual answer
A=3*2.85=8.55
Difference is
D=J-A=5.85-8.55=-2.7